# Emotions Where Art Thou: Understanding and Characterizing the Emotional Latent Space of Large Language Models

### v1

@@ Appendix D ML-AURA Across Models
The ML-AURA analysis in Section 5 was reproduced with Llama3.1-8B-Instruct, Olmov2-Base OLMo et al. (2024) (12), Olmov2-Instruct (12), and Ministral-8B AI (2024) (13) showing results consist with what was found with Llama-3.1. Olmov2-Base performed worse than Olmov2-Instruct. Both of those models performed worse than Ministral-8B and the Llama at the same stage of training. The neurons in the Ministral-8B model were only slightly able than the neurons in the Llama model to separate between emotions. However, in all models more than the majority of neurons at all layers are able to do 1vAll classification of the specified emotion of interest at an AUROC > 0.9 showing great separability in how the different emotions are represented with a low amount of confusion.
@@
Appendix E Emotional Consistency Across Principal Axes
@@
Appendix F Ablations for Emotional Steering
In Section 6, we introduced a method for steering how LLMs internally represent and perceive emotion. This appendix presents ablation studies identifying which components are essential for successful steering. We evaluate the impact of: (1) the number of steering dimensions in the SVD subspace, (2) the presence of the GELU nonlinearity, (3) the use of synonyms in the loss function, (4) the weight of the target-token term in the cross-entropy loss, (5) individual components of the semantic similarity loss, (6) the structure of the margin loss, and (7) the choice of target layers for intervention.
@@ Appendix F Ablations for Emotional Steering
Tables 9 and 10 examines the effect of varying the margin weights $m_{1}$ and $m_{2}$, which define separation constraints in the semantic loss. The margin $m_{1}$ enforces a minimum distance between the target emotion token and its synonyms, preventing collapse and encouraging meaningful local structure. We observe that performance remains relatively stable across $m_{1}$ values, though some instability appears for fear, suggesting mild sensitivity. In contrast, $m_{2}$ enforces separation between the target emotion token and all other emotion tokens (and their synonyms). Steering is highly sensitive to this margin: low $m_{2}$ values consistently fail, while performance improves monotonically as $m_{2}$ increases. At $m_{2}=20$, all emotion-dataset pairs steer successfully, indicating that strong inter-class separation is essential. We adopt $m_{1}=0.75$, $m_{2}=20$ as the best-performing configuration.

### v2

@@ Appendix F ML-AURA Across Models
The ML-AURA analysis in Section 4.3 was reproduced with Llama3.1-8B-Instruct, Olmov2-Base OLMo et al. (2024) (12), Olmov2-Instruct (12), and Ministral-8B (AI, 2024) (13) showing results consist with what was found with Llama-3.1. Olmov2-Base performed worse than Olmov2-Instruct. Both of those models performed worse than Ministral-8B and the Llama at the same stage of training. The neurons in the Ministral-8B model were only slightly able than the neurons in the Llama model to separate between emotions. However, in all models more than the majority of neurons at all layers are able to do 1vAll classification of the specified emotion of interest at an AUROC > 0.9 showing great separability in how the different emotions are represented with a low amount of confusion.
@@
Appendix G Emotional Consistency Across Principal Axes
@@
Appendix H Emotional Steering Per Dataset Results
In Section 4.4, the results of the steering method were presented as averages across English and non-English datasets. In this appendix, the corresponding per-dataset results are presented. Table 8 shows that the learned steering approach achieves consistent and accurate control over internal emotional representations across a diverse set of languages and datasets for all three models. For most emotions, post-steering accuracy typically exceeds 85% across models. Performance is robust even in multilingual settings, with particularly strong results in French, German, and Italian. Steerability remains high for many emotions in Hindi, a lower-resource language, suggesting that lexical sparsity and data imbalance remain limiting factors for certain emotions in under-resourced settings across models. The semantic similarity loss across all datasets, models, and emotions is low, indicating that steering preserves much of the original representational structure while enabling control of emotional perception. LLaMA-3.1-8B is the most steerable model for a plurality of emotions and datasets; however, Olmov2 shows the greatest average delta across emotions and datasets—an outcome notable given that Olmov2-Instruct was previously shown in Section 4 to struggle with representing emotions in a unified manner. Appendix I reports ablations isolating the method’s key factors.
Llama3.1-8B
Ministral
Olmov2
Appendix I Ablations for Emotional Steering
In Section 4.4, we introduced a method for steering how LLMs internally represent and perceive emotion. This appendix presents ablation studies identifying which components are essential for successful steering. We evaluate the impact of: (1) the number of steering dimensions in the SVD subspace, (2) the presence of the GELU nonlinearity, (3) the use of synonyms in the loss function, (4) the weight of the target-token term in the cross-entropy loss, (5) individual components of the semantic similarity loss, (6) the structure of the margin loss, and (7) the choice of target layers for intervention.
@@ Appendix I Ablations for Emotional Steering
Tables 10 and 11 examines the effect of varying the margin weights $m_{1}$ and $m_{2}$, which define separation constraints in the semantic loss. The margin $m_{1}$ enforces a minimum distance between the target emotion token and its synonyms, preventing collapse and encouraging meaningful local structure. We observe that performance remains relatively stable across $m_{1}$ values, though some instability appears for fear, suggesting mild sensitivity. In contrast, $m_{2}$ enforces separation between the target emotion token and all other emotion tokens (and their synonyms). Steering is highly sensitive to this margin: low $m_{2}$ values consistently fail, while performance improves monotonically as $m_{2}$ increases. At $m_{2}=20$, all emotion-dataset pairs steer successfully, indicating that strong inter-class separation is essential. We adopt $m_{1}=0.75$, $m_{2}=20$ as the best-performing configuration.